# Towards Human-AI Collaboration in Healthcare: Guided Deferral Systems with Large Language Models

### v1

@@ Abstract
Large language models (LLMs) present a valuable technology for various applications in healthcare, but their tendency to hallucinate introduces unacceptable uncertainty in critical decision-making situations. Human-AI collaboration (HAIC) can mitigate this uncertainty by combining human and AI strengths for better outcomes. This paper presents a novel guided deferral system that provides intelligent guidance when AI defers cases to human decision-makers. We leverage LLMs’ verbalisation capabilities and internal states to create this system, demonstrating that fine-tuning smaller LLMs with data from larger models enhances performance while maintaining computational efficiency. A pilot study showcases the effectiveness of our deferral system.
@@ 1 Introduction
Clinicians often make decisions using their expertise in addition to the intelligent guidance of colleagues, which we define as task-based recommendations and informed reasoning rooted in logic. Current deferral systems lack this guidance, isolating human decision-makers. We propose that effective deferral systems should also provide intelligent guidance. This paper explores using LLMs to achieve this goal.
The main challenge in building such systems is the computational expense of LLMs. Proprietary LLMs offer advantages such as state-of-the-art performance and easy implementation without needing high-performance hardware. However, they are impractical for data-sensitive applications due to lack of internal state access and privacy concerns. Open-source LLMs can perform well with relatively large amounts of parameters, but are slow and require high-performance hardware. Smaller LLMs are faster but less effective. These issues affect all real-world LLM applications, not just deferral systems. In this paper, we show how instruction-tuning small, efficient open-source LLMs with data from larger models can improve performance, maintain efficiency on consumer hardware, and ensure data privacy.
@@
  • We demonstrate that instruction-tuning an open-source, efficient and small-scale LLM on the guard-railed generation of a large-scale version of the same LLM leads to a much improved classification performance and deferral performance, surpassing even that of the latter.
@@ 2 Related Work
Few studies explore the use of LLMs in HAIC. Wiegreffe et al. (2022) first examined LLMs for explaining classification decisions using a human-in-the-loop approach to train a filter assessing explanation quality. Rastogi et al. (2023) used HAIC to audit error-prone LLMs with other LLMs. Dvijotham et al. (2023) proposed CoDoC, a deferral system with a black-box classifier and learned deferral AI for healthcare. Our system differs by using LLMs for guided deferrals. Banerjee et al. (2023) used LLMs to provide textual guidance for clinicians in decision-making on clinical imaging tasks and argued that deferral systems are sub-optimal due to anchoring bias and lack of decision-maker support. They proposed learning-to-guide, which trains LLMs to provide decision-making guidance without deferring cases. We contend that this approach still burdens decision-makers with the time and fatigue issues deferral systems address. Our work combines LLMs in deferral systems with valuable guidance for decision-makers on deferred cases, laying the foundation for advanced deferral algorithms like Learning-to-Defer (Madras et al., 2018) or Learning with Rejection (Cortes et al., 2016). Few studies (Mozannar & Sontag, 2021; Liu et al., 2021) have evaluated the efficacy of HAIC on text classification tasks but are not LLM focused.
Existing research on LLMs in selective prediction includes methods to measure uncertainty in model responses after generation (Varshney & Baral, 2023). Chen et al. (2023) proposed improving selective prediction by incorporating self-assessment. Ren et al. (2023) explored detecting out-of-distribution instances in summarization and translation tasks. Our paper applies selective prediction to deferral systems in clinical parsing. We show that combining the model’s internal state with its generated prediction enhances selective prediction without post-generation methods. We provide a comprehensive evaluation of selective prediction performance in clinical classification using real-world data, applied in deferral systems with in-distribution data.
@@ 3.1 Sources of Predictions
Hidden-State Prediction Source. The second prediction probability is defined based on the hidden representations of LLM to implicitly detect classifications. This is inspired by Ren et al. (2023), which utilises the final-layer hidden embedding of LLMs for out-of-distribution detection. The output embedding for input $\boldsymbol{x}_{i}$, $\hat{h}_{i}$, was computed as the average of the decoder’s final-layer hidden-state vectors $\boldsymbol{g}_{ik}\in\mathbb{R}^{d}$ over all $K$ output tokens with a hidden dimension of $5120$ for the small-scale LLM:
@@
A 3-layer MLP is then trained as a hidden-state classifier to learn the probability of having the disorder from the LLM hidden representation, the hidden-state prediction probabilities, denoted $\hat{\epsilon}_{i}$. We experimented with different models for the hidden-state classifier in ambitions of fully utilising the information embedded in $\hat{h}$ (see Appendix D.1 for details).
@@ 3.2 Instruction-Tuning Methodology
In generating well-calibrated verbalised probabilities, we use the “verb. 1S top-$k$” prompting strategy (Tian et al., 2023), which prompts the LLM to provide the top $k$ guesses and their probabilities in a single response. Adapting this to $k=1$ for a binary setting, we prompt the LLM to return the probability of the positive class. Additionally, we ask for the top reason the disorder may be present and the top reason it may not be present, using dialectic reasoning (Hegel, 2018) to provide intelligent guidance for decision-makers. This technique has proven effective in decision support (Jarupathirun & Zahedi, 2007). An example output is shown in Listing LABEL:lst:instruction_output.
Instruction-Tuning Data Generation. We use 4-bit quantized 13B and 70B versions of the open-source LLM Tulu V2 DPO (Ivison et al., 2023), a Llama 2 derivative (Touvron et al., 2023), as our small and large-scale LLMs. As classification and deferral performance improved with model size (see Table 1), we instruction-tuned the small-scale model using guard-railed answers generated by the large-scale model. We found that tuning with 16 generated answers per example yielded the best validation set deferral performance. We implement guard-rails to reject answers containing hallucinations or “bad” behaviours, such as non-English language, lack of conciseness, or illogical statements. Details on guard-rails are in Appendix B.3, and the base prompt used is in Appendix A Listing LABEL:lst:instruction. We instruction-tune the small-scale LLM using QLoRA (Dettmers et al., 2023). Appendix B contains experimental details for reproducibility.
TOP REASON FOR:
The MRI report indicates that there is narrowing of the right exit foramen at the L5-S1 level, which is causing compression of the exiting right L5 nerve root. This suggests that there is foraminal stenosis (FS) at this level.
TOP REASON AGAINST:
There is no mention of foraminal stenosis specifically at the L5-S1 level.
CONCLUSION:
Based on the information provided, there is a possibility of foraminal stenosis at the L5-S1 level due to the narrowing of the right exit foramen and compression of the exiting right L5 nerve root. However, the report does not explicitly mention foraminal stenosis at this level.
@@ CONCLUSION:
Listing 1: Example guidance output based on a spinal MRI radiological report. The instruction-tuned LLM is able to intelligently infer its presence with sound logic even without explicit diagnosis of a disease.
@@ 3.3 Deferral Mechanism
The deferral performance is measured through recursively iterating through all elements $\tilde{p}$, deferring this prediction (without replacement) and measuring the classification performance of the LLM on the remaining cases. Intuitively, good deferral behaviour should demonstrate a monotonically increasing accuracy with increasing deferral rate.
@@ 3.4 Pilot Study in Investigating the Effectiveness of Generated Guidance
We conducted a pilot study with 20 participants to evaluate the effectiveness of our guidance when deferring 30 ($\approx$5%) of the most uncertain test predictions. Participants received background information, including clinical details and examples of prediction outcomes with associated MRI reports. We also presented the LLM’s performance on a validation set to help participants develop a mental model of the LLM’s capabilities. The 30 questions were randomised for each participant to reduce order bias.
Participants moved to the next question if their prediction matched the LLM’s. If it differed, they received guidance and could either change their prediction or keep it based on the guidance and their understanding of the LLM. This process allowed us to assess human performance with and without guidance. Effective guidance should help participants recognise the accuracy of their judgements. Full details are in Appendix C.
@@
Setup Clf. Perf. Calibration Perf. Deferral Perf. LLM Efficiency
@@
(1)    $\hat{t}_{\text{BASE-13B}}$ 0.61±0.01 0.26±0.01 0.32±0.01 0.897460±0.0063 |||| |||| ||||
(2)    $\hat{\epsilon}_{\text{BASE-13B}}$ 0.80±0.02 0.03±0.01 0.04±0.01 0.994715±0.0006 0.08±0.02 6.02 0.29
(3)    $\hat{\mu}_{\text{BASE-13B}}$ 0.82±0.03 0.13±0.01 0.18±0.01 0.993318±0.0010 |||| |||| ||||
(4)    $\hat{t}_{\text{BASE-70B}}$ 0.83±0.01 0.27±0.01 0.33±0.01 0.973529±0.0052 |||| |||| ||||
(5)    $\hat{\epsilon}_{\text{BASE-70B}}$ 0.91±0.01 0.02±0.01 0.05±0.01 0.997191±0.0002 1.00 32.08 0.04
(6)    $\hat{\mu}_{\text{BASE-70B}}$ 0.92±0.01 0.14±0.01 0.19±0.01 0.997260±0.0003 |||| |||| ||||
(7)    $\hat{t}_{\text{INSTRUCT-13B}}$ 0.89±0.01 0.27±0.01 0.32±0.01 0.987501±0.0024 |||| |||| ||||
(8)    $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ 0.93±0.01 0.01±0.01 0.05±0.01 0.997409±0.0002 0.08±0.02 6.02 0.0002
(9)    $\hat{\mu}_{\text{INSTRUCT-13B}}$ 0.94±0.01 0.14±0.01 0.19±0.01 0.996978±0.0013 |||| |||| ||||
@@ 4 Experiments
We use the OSCLMRIC (Oxford Secondary Care Lumbar MRI Cohort) dataset, containing professionally annotated lumbar MRIs and radiological reports for various types of stenosis at different spinal levels, of which serve as our ground-truth labels. An example report is shown in Listing LABEL:lst:example_report. The dataset is highly imbalanced, with about 95% of labels negative. Each report is parsed to detect the binary presence of three types of stenosis (foraminal stenosis [FS], spinal canal stenosis [SCS], and lateral recess stenosis [LRS]) at six lumbar spine levels, resulting in 1,800 examples. The data is randomly split into 30% (540) for generating an instruction-tuning dataset, 20% (360) for training the hidden-state classifier, 20% (360) for validation, and 30% (540) for testing.
SOTA Baseline. We use Tulu V2 70B as the SOTA baseline model for our experiments, of which is the highest-performing open-source LLM against several benchmarks at the time these experiments were conducted (Beeching et al., 2023). Due to the aforementioned data privacy issues with proprietary LLMs, we omit these models as baselines.
Abbreviations. We abbreviate instruction-tuned and base (non instruction-tuned) models as “INSTRUCT” and “BASE”, respectively. Tests for statistical significance (SS) are conducted using a one-sided non-parametric Mann-Whitney U test (Mann & Whitney, 1947).
Evaluation Metrics. The Area Under Accuracy-Rejection Curve (AUARC) (Nadeem et al., 2009) and F1-score are used to measure the deferral and classification performance, respectively. We further evaluate the calibration of our setups with ECE (Guo et al., 2017) and ACE Nixon et al. (2020). LLM efficiency is evaluated based on Rel. s/Gen. (seconds/generation relative to the baseline), Mem. (GPU VRAM required in GB), and E.R. the error rate (the number of unsuccessful generations for each successful generation).
@@ 4.1 Computational Results
We find a SS improvement in F1-Score of verbalised predictions of $\hat{t}_{\text{INSTRUCT-13B}}$ in comparison to the baseline $\hat{t}_{\text{BASE-70B}}$, demonstrating instruction-tuning methodology to be successful. The greatest deferral performance results from the hidden-state predictions of the 13B setup, surpassing that of the 70B baseline. The instruction-tuned model is seen to significantly improve in all LLM efficiency metrics.
@@
The Combined Prediction Leads to Greatest Classification Performance. We combine the verbalised prediction and the hidden state prediction into a single prediction simply through their mean, of which we denote $\hat{\mu}$. $\hat{\mu}_{\text{INSTRUCT-13B}}$ is the greatest classification source (Table 1, Row 9), surpassing that of $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ and $\hat{t}_{\text{INSTRUCT-13B}}$. We find a SS improvement over all other sources ($p<$0.01). This implies an interesting finding, in that the verbalised and hidden-state sources contain valuable and distinct pieces of information contributing to classification. We theorise that this is because $\hat{t}$ acts as a regulariser on $\hat{\epsilon}$. See Appendix D.2 for additional analysis.

### v2

@@ Abstract
Large language models (LLMs) present a valuable technology for various applications in healthcare, but their tendency to hallucinate introduces unacceptable uncertainty in critical decision-making situations. Human-AI collaboration (HAIC) can mitigate this uncertainty by combining human and AI strengths for better outcomes. This paper presents a novel guided deferral system that provides intelligent guidance when AI defers cases to human decision-makers. We leverage LLMs’ verbalisation capabilities and internal states to create this system, demonstrating that fine-tuning small-scale LLMs with data from large-scale LLMs greatly enhances performance while maintaining computational efficiency and data privacy. A pilot study showcases the effectiveness of our proposed deferral system.
@@ 1 Introduction
Clinicians often make decisions using their expertise in addition to the intelligent guidance of colleagues, which we define as task-based recommendations and informed reasoning rooted in logic. Current deferral systems lack this guidance, isolating human decision-makers. We propose that effective deferral systems should additionally simulate providing this intelligent guidance to decision-makers. This paper explores using LLMs to achieve this.
The main challenge in building such systems is the computational expense of LLMs. Proprietary LLMs offer advantages such as state-of-the-art performance and easy implementation without needing high-performance hardware. However, they are impractical for data-sensitive applications due to lack of internal state access and privacy concerns. Open-source LLMs can perform well with relatively large amounts of parameters, but are slow and require high-performance hardware. Smaller LLMs are faster but less effective. These issues affect all real-world LLM applications, not just deferral systems. In this paper, we propose methodology for developing efficient and accurate LLMs of which are capable of guided deferral and suitable for healthcare applications.
@@
  • We demonstrate that instruction-tuning an open-source, efficient, and small-scale LLM on the guard-railed generation of a large-scale version of the same LLM leads to significantly improved classification and deferral performance in this task, surpassing even that of the larger model through the use of guardrails. This methodology ensures data privacy, which is critical for high-stakes environments such as healthcare.
@@ 2 Related Work
Human-AI Collaboration with Large Language Models. Few studies explore the use of LLMs in HAIC. Wiegreffe et al. (2022) first examined LLMs for explaining classification decisions using a human-in-the-loop approach to train a filter assessing explanation quality. Rastogi et al. (2023) used HAIC to audit error-prone LLMs with other LLMs. Dvijotham et al. (2023) proposed CoDoC, a deferral system with a black-box classifier and learned deferral AI for healthcare. Our system differs by using LLMs for guided deferrals. Banerjee et al. (2023) used LLMs to provide textual guidance for clinicians in decision-making on clinical imaging tasks and argued that deferral systems are sub-optimal due to anchoring bias and lack of decision-maker support. They proposed learning-to-guide, which trains LLMs to provide decision-making guidance without deferring cases. We contend that this approach still burdens decision-makers with the time and fatigue issues deferral systems address. Our work combines LLMs in deferral systems with valuable guidance for decision-makers on deferred cases, laying the foundation for advanced deferral algorithms like Learning-to-Defer (Madras et al., 2018) or Learning with Rejection (Cortes et al., 2016). Few studies (Mozannar & Sontag, 2021; Liu et al., 2021) have evaluated the efficacy of HAIC on text classification tasks but are not LLM focused.
Selective Prediction of Large Language Models. Existing research on LLMs in selective prediction includes methods to measure uncertainty in model responses after generation (Varshney & Baral, 2023). Chen et al. (2023) proposed improving selective prediction by incorporating self-assessment. Ren et al. (2023) explored detecting out-of-distribution instances in summarisation and translation tasks. Our paper applies selective prediction to deferral systems in clinical parsing. We show that combining the model’s internal state with its generated prediction enhances selective prediction without post-generation methods. We provide a comprehensive evaluation of selective prediction performance in clinical classification using real-world data, applied in deferral systems with in-distribution data.
Instruction-Tuning of Large Language Models. Finally, whilst there exists literature demonstrating improved zero-shot performance of instruction-tuned (IT) LLMs on unseen tasks (Wei et al., 2022), our work studies this improvement specifically on an in-domain task through the use of guardrails. Additionally, there exists works on the use of IT in the medical domain (Liu et al., 2023), but these focus on report summarisation. No works have researched the applications of IT LLMs for the use of deferral systems. Zhang et al. (2023) provide a recent survey for additional insight in this area.
@@ 3.1 Sources of Predictions
Hidden-State Prediction Source. The second prediction probability is defined based on the hidden representations of LLM to implicitly detect classifications. This is inspired by Ren et al. (2023), which utilises the final-layer hidden embedding of LLMs for out-of-distribution detection. The output embedding $\hat{h}_{i}\in\mathcal{H}$ for input $\boldsymbol{x}_{i}$ is computed as the average of the decoder’s final-layer hidden-state vectors $\boldsymbol{g}_{ik}\in\mathbb{R}^{d}$ over all $K$ output tokens with a hidden dimension of $d=5120$ for the small-scale LLM:
@@
Then, an MLP $g:\mathcal{H}\mapsto\mathbb{R}$ is trained as a hidden-state classifier to learn the probability of the disorder from the LLM hidden representation, the hidden-state prediction probabilities, denoted $\hat{\epsilon}_{i}$, i.e. $\hat{\epsilon}_{i}=g(\hat{h}_{i})$. We experimented with different models for the hidden-state classifier in ambitions of fully utilising the information embedded in $\hat{h}$ (see Appendix D.1 for details).
Combined Prediction. Additionally, we combine the verbalised prediction and the hidden state prediction through their mean to form the combined prediction, of which we denote $\hat{\mu}_{i}$.
@@ 3.2 Instruction-Tuning Methodology
In generating well-calibrated verbalised probabilities, we use the “verb. 1S top-$k$” prompting strategy (Tian et al., 2023), prompting the LLM to provide the top $k$ guesses and their probabilities in a single response. Adapting this to $k=1$ for a binary setting, we prompt the LLM to return the probability of the positive class. Additionally, we prompt for the top reason the disorder might and might not be present, using dialectic reasoning (Hegel, 2018) to provide intelligent guidance for decision-makers. This technique has proven effective in decision support (Jarupathirun & Zahedi, 2007). An example output is in Figure 2.
Example Generated Guidance TOP REASON FOR: The MRI report indicates that there is narrowing of the right exit foramen at the L5-S1 level, which is causing compression of the exiting right L5 nerve root. This suggests that there is foraminal stenosis (FS) at this level.
TOP REASON AGAINST: There is no mention of foraminal stenosis specifically at the L5-S1 level.
CONCLUSION: Based on the information provided, there is a possibility of foraminal stenosis at the L5-S1 level due to the narrowing of the right exit foramen and compression of the exiting right L5 nerve root. However, the report does not explicitly mention foraminal stenosis at this level.
@@
Figure 2: Example guidance based on a spinal MRI radiological report. The instruction-tuned LLM is able to intelligently infer a diagnosis with sound logic without explicit textual diagnosis.
@@ 3.3 Deferral Mechanism
The deferral performance is measured through recursively iterating through all elements $\tilde{p}$, deferring this prediction (without replacement) and measuring the classification performance of the LLM on the remaining cases. This procedure describes the AUARC metric (Nadeem et al., 2009) when measuring accuracy. Intuitively, good deferral behaviour should demonstrate a monotonically increasing accuracy with increasing deferral rate.
@@ 3.4 Pilot Study in Investigating the Effectiveness of Generated Guidance
We conduct a pilot study with 20 participants to evaluate the effectiveness of our guidance in the scenario of deferring 30 ($\approx$5%) of the most uncertain test predictions. Participants received background information, including clinical details, examples of prediction outcomes with associated MRI reports and the LLM’s performance on a validation set to help participants develop a sufficient mental model of the LLM, of which has been shown to be important in HAIC systems (Kulesza et al., 2012; Bansal et al., 2019). The ordering of the 30 questions were randomised for each participant to reduce order bias.
Participants moved to the next question if their prediction matched the LLM’s. If it differed, they received guidance and could either change their prediction or keep it based on the guidance and their understanding of the LLM. This process allowed us to assess human performance with and without guidance. Effective guidance should help participants recognise both the accuracy and inaccuracy of their judgements. Full details are in Appendix C.
@@
Setup Classification Perf. Calibration Perf. Deferral Perf. LLM Efficiency
@@
(1)    $\hat{t}_{\text{BASE-13B}}$ 0.85±0.02 0.61±0.01 0.61±0.01 0.26±0.01 0.32±0.01 0.897460±0.0063 |||| |||| ||||
(2)    $\hat{\epsilon}_{\text{BASE-13B}}$ 0.73±0.02 0.94±0.03 0.80±0.02 0.03±0.01 0.04±0.01 0.994715±0.0006 0.08±0.02 6.02 0.29
(3)    $\hat{\mu}_{\text{BASE-13B}}$ 0.78±0.03 0.87±0.04 0.82±0.03 0.13±0.01 0.18±0.01 0.993318±0.0010 |||| |||| ||||
(4)    $\hat{t}_{\text{BASE-70B}}$ 0.90±0.02 0.78±0.01 0.83±0.01 0.27±0.01 0.33±0.01 0.973529±0.0052 |||| |||| ||||
(5)    $\hat{\epsilon}_{\text{BASE-70B}}$ 0.87±0.01 0.97±0.01 0.91±0.01 0.02±0.01 0.05±0.01 0.997191±0.0002 1.00 32.08 0.04
(6)    $\hat{\mu}_{\text{BASE-70B}}$ 0.90±0.01 0.95±0.02 0.92±0.01 0.14±0.01 0.19±0.01 0.997260±0.0003 |||| |||| ||||
(7)    $\hat{t}_{\text{INSTRUCT-13B}}$ 0.85±0.02 0.92±0.02 0.88±0.02 0.27±0.01 0.32±0.01 0.987501±0.0024 |||| |||| ||||
(8)    $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ 0.89±0.01 0.98±0.01 0.93±0.01 0.01±0.01 0.05±0.01 0.997409±0.0002 0.08±0.02 6.02 0.0002
(9)    $\hat{\mu}_{\text{INSTRUCT-13B}}$ 0.91±0.01 0.98±0.01 0.94±0.01 0.14±0.01 0.19±0.01 0.996978±0.0013 |||| |||| ||||
@@ 4 Experiments
We use the OSCLMRIC (Oxford Secondary Care Lumbar MRI Cohort) dataset, containing professionally annotated lumbar MRIs and radiological reports for various types of stenosis at different spinal levels, of which serve as our ground-truth labels. An example report is shown in Listing LABEL:lst:example_report. The dataset is highly imbalanced, with $\approx$95of labels negative. Each report is parsed to detect the binary presence of three types of spinal stenosis (foraminal stenosis [FS], spinal canal stenosis [SCS], and lateral recess stenosis [LRS]) at six lumbar spine levels, resulting in 1,800 examples. The data is randomly split into 30% for generating an instruction-tuning dataset, 20% for training the hidden-state classifier, 20% for validation, and 30% for testing.
SOTA Baseline. We use Tulu V2 70B as the SOTA baseline model for our experiments, of which is the highest-performing open-source LLM against several benchmarks at the time these experiments were conducted (Beeching et al., 2023). Proprietary LLMs are omitted as baselines due to the aforementioned data privacy issues.
Abbreviations. We abbreviate IT and base (non-IT) models as “INSTRUCT” and “BASE”, respectively. Tests for statistical significance (SS) are conducted using a one-sided non-parametric Mann-Whitney U test (Mann & Whitney, 1947), unless stated otherwise.
Evaluation Metrics. The Area Under Accuracy-Rejection Curve (AUARC) and F1-score are used to measure the deferral and classification performance, respectively. We further evaluate the calibration of our setups with ECE (Guo et al., 2017) and ACE (Nixon et al., 2020). LLM efficiency is evaluated based on Rel. s/Gen. (seconds/generation relative to the baseline), Mem. (GPU VRAM required in GB), and E.R. the error rate (the number of unsuccessful generations for each successful generation).
@@ 4.1 Computational Results
We find a SS improvement in F1-Score of verbalised predictions of $\hat{t}_{\text{INSTRUCT-13B}}$ in comparison to the baseline $\hat{t}_{\text{BASE-70B}}$, demonstrating instruction-tuning methodology to be successful. The greatest deferral performance results from the hidden-state predictions of the 13B setup, surpassing that of the 70B baseline. The IT model is seen to significantly improve in all LLM efficiency metrics.
@@
The Combined Prediction Leads to Greatest Classification Performance. $\hat{\mu}_{\text{INSTRUCT-13B}}$ is the greatest classification source (Table 1, Row 9), surpassing that of $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ and $\hat{t}_{\text{INSTRUCT-13B}}$ ($p$$<$0.01). The improvement in $\hat{\mu}$ over $\hat{t}$ and $\hat{\epsilon}$ is seen in all setups. This implies an interesting finding; in that the verbalised and hidden-state sources contain valuable and distinct pieces of information contributing to classification. There exists a positive correlation of $\rho=0.53$ $0.07$ between these predictions. When focusing our analysis when $\hat{\mu}$ is correct and the predictions of $\hat{t}$ and $\hat{\epsilon}$ differ, a correct $\hat{\epsilon}$ prediction is combining with an incorrect $\hat{t}$ prediction in the majority of cases, at an average of (75 ± 0.14)of these cases. When $\hat{\mu}$ is incorrect, a correct $\hat{t}$ prediction is combining with an incorrect $\hat{\epsilon}$ prediction in the majority of cases, at an average of (75 ± 0.25)of these cases. See Figure 3 for confusion matrices.
Figure 3: Confusion matrix illustrating the number of test cases for $\hat{t}$ and $\hat{\epsilon}$ precision, given (i) the incorrect prediction of their combination $\hat{\mu}$, and (ii) correct prediction.
The Combined Prediction Has The Greatest Deferral Performance in Mostly-Autonomous Systems. We find a SS ($p$$<$0.1) improvement in the deferral strategy of deferring on $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ and classifying on $\hat{\mu}_{\text{INSTRUCT-13B}}$, in comparison to the next-best strategy of deferring and classifying solely on $\hat{\epsilon}_{\text{INSTRUCT-13B}}$ , in situations where $<$5% of cases are deferred. In practice, this would be cases where users would require the system to be mostly autonomous. This implies that the combined prediction is a greater predictor of disease in high-uncertainty situations. In deferral rates $\geq$5%, there exists no SS difference between the two strategies, but a SS over all other setups.